# Automated test generation to evaluate tool-augmented LLMs as conversational AI agents

### v1

@@ 1 Introduction
  • ALMITA, the first conversational dataset that can be used to evaluate customer support AI agents, including both tooling (i.e., functions) and conversation reply to follow company user-defined procedures. ALMITA contains 1420 synthetic tests that were manually curated to ensure high-quality samples [^1] ALMITA, along with all other datasets generated using our pipeline and referenced in the paper, will be made available upon acceptance of the paper..
@@ 3.8 Conversation generator
The conversation generator creates synthetic conversations from an input conversation graph, a sampled path from the conversation graph, and relevant APIs. We provide the LLM with context about the conversation graph structure and the APIs. Using one-shot prompting, we present the LLM with an example triplet consisting of a conversation graph, a list of APIs, and a sampled path, as well as a possible conversation based on these conditions (see Section A.6). In an effort to generate valid conversations, we include specific conditions in the prompt, such as always generating a message with the API output following an API message, alternating customer and assistant messages, ensuring agents act on API output messages, and verifying API input and output types.
@@ 4 Results
In Section 4.1, we detail the creation of ALMITA, a curated dataset for evaluating LLMs as AI customer support agents. To ensure dataset quality, we employ manual filtering. Two annotators independently review each datapoint to identify incorrect instances, followed by a discussion to align their assessments and minimize disagreements. Any datapoint deemed incorrect by at least one annotator is then removed. GPT-4 is used for all generation steps (see Figure 1). To assess the benefits of the graph intermediate structures, we conduct an ablation study comparing conversations generated directly from procedures to those using the intermediate structures, with manual curation for quality assessment (Section 4.2). In Section 4.3, we evaluate various AI agents on ALMITA. Finally, in Section 4.4, we assess the effectiveness of our pipeline in generating high-quality test sets automatically. We do this by comparing the AI agents’ performance on ALMITA with those on its fully automated counterpart, auto-ALMITA.

### v2

@@ 1 Introduction
  • ALMITA, the first conversational dataset that can be used to evaluate customer support AI agents, including both tooling (i.e., functions) and conversation reply to follow company user-defined procedures. ALMITA contains 1420 synthetic tests that were manually curated to ensure high-quality samples [^1] ALMITA, along with all other datasets generated using our pipeline and referenced in the paper, are available in https://github.com/zendesk/almita-dataset..
@@ 2 Related work
Other relevant work focuses on graph learning and on using different intermediate structures to reducing hallucinations. ye2023natural propose InstructGLM, which uses natural language to describe node features used to tune an LLM for inference on graphs. wang2024can introduce NLGraph, a benchmark for graph-based problems written in natural language, demonstrating that LLMs can perform structured operations on textual descriptions of graphs. Additionally, narayan2023conditional propose using question-answer blueprints as intermediate representations to reduce hallucinations. These works do not fully encompass our problem setting of generating conversations in dialog format, calling APIs, and extracting tests.
@@ 3.8 Conversation generator
The conversation generator creates synthetic conversations from an input conversation graph, a sampled path, and relevant APIs. We provide the LLM with context about the conversation graph structure and the APIs. Using one-shot prompting, we present the LLM with an example triplet consisting of a conversation graph, a list of APIs, and a sampled path, as well as a possible conversation based on these conditions (see Section A.6). In an effort to generate valid conversations, we include conditions in the prompt, such as always generating a message with the API output following an API message, alternating customer and assistant messages, ensuring agents act on API output messages, and verifying API input and output types.
@@ 4 Results
In Section 4.1 we detail the creation of ALMITA, a manually curated dataset for evaluating LLMs as AI customer support agents. Two annotators independently review each datapoint to identify incorrect instances, followed by a discussion to align their assessments and minimize disagreements. Any datapoint deemed incorrect by at least one annotator is then removed. GPT-4 is used for all generation steps (see Figure 1). To assess the benefits of the graph intermediate structures, we conduct an ablation study comparing conversations generated directly from procedures to those using the intermediate structures, with manual curation for quality assessment (Section 4.2). In Section 4.3, we evaluate various AI agents on ALMITA. Finally, in Section 4.4, we assess the effectiveness of our pipeline in generating high-quality test sets automatically. We do this by comparing the AI agents’ performance on ALMITA with those on its fully automated counterpart, auto-ALMITA.